TOTALLY DISCONNECTED JULIA SETS

Peter Makienko

Abstract. We apply the pinching and plumbing constructions to study the boundaries of spaces of quasiconformal deformations of rational maps with totally disconnected Julia sets.

Typeset by AMS-TEX